Compute the pH of a solution at 25 degree centigrade which is twice as alkaline as pure water

Grade:12

1 Answers

vikas askiitian expert
13 years ago

ph of water = 7

concentration of [OH-] = 10-7

concentration of [OH-] in the solution  = 2*10-7                     (2 times alkaline than water)

total OH- in sol is 10-7  + 2*10-7 =3*10-7

[H+] = 10-14/3*10-7 = 3.33*10-8

PH = -log3.33*10-8 =8 - log3.33 = 7.5 approx